[Brahms-l] Shift report 20040321 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sun Mar 21 2004 - 16:00:24 EST
------------------------------------------------------------
Date: Sun Mar 21 16:00:23 2004, Shift: 08:00-16:00
Supervisor: PS, with: SN
Log book#15 pages 94 - 94

Shift Summary:
------------------------------------------------------------
1. Nice beam most of the shift.

2. 4 deg   tr2    tr3    tr6
   11218   112k   113k   164k
   11219   215k   242k   347k
   11220   220k   219k   336k
   11221   174k   197k   285k
   TOTAL   720k   771k   1131k

   8 deg
   11222   51k    108k   97k
   11223   42k    124k   91k
   11224   40k    154k   112k
   TOTAL   134k   386k   301k

   11225   in progress...

3. Everythng ran smoothly. Completed 4 deg setting and moved
   FS to to 8 deg. 
   Polarity changed on all magnets throughout shift.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 21 16:01:06 2004

This archive was generated by hypermail 2.1.8 : Sun Mar 21 2004 - 16:01:23 EST